How many moles of hydrazine are needed to make 3.12 mol of dinitrogen tetraoxide react?
1
Expert's answer
2018-02-05T07:17:00-0500
.n(N2O4)=3.12 mol
.n(N2H4)=?
The reaction is 2 N2H4+N2O4=3N2+4H2O
n(N2H4)=2 n (N2O4)= 2 *3.12=6.24 mol